OpenStudy (anonymous):

Solve 5(2x + 1) = 12x – 1

OpenStudy (anonymous):

10x + 5 = 12x - 1 10x - 12x = -1-5 -2x= -6 x=3

OpenStudy (anonymous):

X = 3

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

thanks for giving me medal

OpenStudy (anonymous):

:( no medal for second man in a?

OpenStudy (anonymous):

YA! thank you

OpenStudy (anonymous):

your welcome! :)